if cos A= 8/10 , which of the following is true?
A) Tan A= 4/3
B) Sin A = 4/5
C) Tan A= 3/4
D) Sin A = 5/3

Answer by Cromlix(4381)

Hi there,
If cosA = 8/10
then sinA = 6/10
TanA = 6/8 or 3/4
Answer C)
Hope this helps :-)